[figure: 9.7 cm]

HUBBARD MODEL APPROACH TO THE ELECTRON TRANSPORT IN A MOLECULAR DEVICE

\underline{Kamil Walczak} and Tomasz Kostyrko Institute of Physics, A.Mickiewicz University Umultowska 85, 61 - 614 Poznań, Poland

Electronic transport through a molecular device including benzene-like molecule connected to two semi-infinite leads is studied using the scattering approach. The molecule is described with the help of the single-orbital π-electron approximation with the electron interactions treated within the Hubbard model. The D-dimensional hypercubic leads (D=1, 2, 3) are described using the single band tight-binding model. The electron interactions in the molecule are treated by means of the Hartree-Fock approximation and the effects of selfconsistency on the transport properties of the device are analyzed in detail. The transmission function for a finite bias voltage is computed and the results are exploited to obtain I-V characteristics of the device. It is shown that the transmission is determined by the combined effect of the electronic structure of the molecule and the molecule-to-lead coupling. The Hubbard interaction neutralizes voltage-induced charge redistribution in the molecule which results in a shift of the transmission resonances and a substantial change of the current for intermediate values of voltage.

Work supported in part by K.B.N. Poland, project number 2 PO3B 087 19.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Kamil Walczak

Address for correspondence : Institute of Physics, A.Mickiewicz University Umultowska 85, 61 - 614 Poznań, Poland

Email address : walczak@amu.edu.pl